OpenStudy (anonymous):

Having a square bedroom with 5 mtrs as the length of the diagonal, compute for the number of 300mmx 300mm floor tile.. can you answer this question? :) thnx!!~

OpenStudy (anonymous):

area of the room is 5*5/2=25/2 sqm hence the number of tiles reqd is 25*1000*1000/(2*300*300)=2500/18=139

OpenStudy (anonymous):

convert 300mm to meters-> 300mm x 1m/1000mm= 0.3meters. so you have 0.3mx0.3m floor tiles If the diagonal of the room is 5 meters, and the room is a square, then you can compute the 2 side lengths using Pythagoras: a\[a ^{2} + b ^{2} = 5^{2}\] since a = b, \[2a ^{2}=25\] Solve for a, and you get the side length of the room to be \[\sqrt{12.5}\] so the area is the side length squared divided by 2, witch is 6.25m squared. If each floor tile is 0.3mx0.3m, it is 0.09m squared. You then divide the area by the tile area, \[6.25\div0.09 = 69.444.. \]